# Customizing Spider Silk: Generative Models with Mechanical Property Conditioning for Protein Engineering

### v1

@@ 3.1 Stage 1: Distillation of ProtGPT2 on Spider Sequences
ProtGPT2 [20]stands out among generative protein language models (PLMs) due to its specialized training and demonstrated ability to generate biologically plausible protein sequences. Unlike large-scale PLMs such as ESM-2 [40] and ProteinMPNN [41], which focus on structure prediction and sequence design constrained by existing protein scaffolds, ProtGPT2 is designed specifically for de novo protein generation. It has been trained on 10 million protein sequences from the UniRef50 dataset of the UniProt Knowledgebase (UniProtKB) [21] using a self-supervised learning approach. This extensive pre-training enables ProtGPT2 to predict the next amino acid in a sequence, effectively capturing the underlying "grammar" of protein structures.
@@
To address this challenge, we employ a multilevel strategy that leverages a pre-trained protein language model while optimizing it for our specific task. We begin with ProtGPT2. To enhance efficiency and adapt the model for MaSp repeat generation, we apply knowledge distillation [42], creating a smaller, task-specific variant: SpiderGPT. This distilled model retains the essential knowledge of its teacher while significantly reducing model size and improving inference time, making it more practical for generating novel MaSp repeat sequences tailored to specific mechanical properties.
The distillation follows a teacher-student framework:
  • Teacher: ProtGPT2, the pre-trained model, serves as the knowledge source.
  • Student: SpiderGPT, a lightweight model trained to replicate the teacher’s outputs.
Dataset: SpiderGPT was trained on a curated dataset of approximately 100k protein sequences obtained from UniProtKB. These sequences were specifically selected based on their taxonomic classification within Araneae (spiders) and an annotation score greater than 1. (Details of the procedure are provided in Appendix A.1.) This focused dataset enables the student model to specialize in spider proteins while leveraging the broader knowledge distilled from the teacher model.
Training Process: The distillation parameters include Temperature $(T)$ to be 10, Alpha $(\alpha)$ to be 0.1, embedding dimension $(n_{embd})$ to be 512, number of transformer layers $(n_{layer})$ to be 6, number of attention heads $(n_{heads})$ to be 8. This reduced architecture was designed to strike a balance between computational efficiency and representational capacity, ensuring that the student model effectively captures domain-specific patterns. To facilitate the distillation process, both the teacher and student models were trained on the same tokenized dataset. The teacher model generated probability distributions over the vocabulary for each input sequence, serving as soft labels for the student model.
@@ 3.2 Stage 2: Level 1 Fine-Tuning on MaSp Repeat Regions
Dataset: For this study, we utilized the Spider Silkome dataset [15], a comprehensive resource cataloging silk gene sequences from 1,098 spider species and measuring mechanical, thermal, structural, and hydration properties for 446 species. It highlights the role of MaSp paralogs (MaSp1–MaSp3) in high-performance silk and identifies key amino acid motifs contributing to silk properties. This dataset serves as an open platform for advancing biomaterial research and innovation.
@@
For level 1 fine-tuning, we curated a dataset of MaSp sequences, resulting in 6,000 instances. To focus on the functional repeat regions, we removed the first 150 and last 115 amino acid residues, which correspond to the N-terminal and C-terminal domains, respectively [43]. The repeat sequences of the same MaSp type were then concatenated for each species to create a structured dataset.
Training Process: The fine-tuning process employed a causal language modeling objective, which fundamentally leverages the transformer architecture’s auto-regressive nature. This approach ensures that the model learns to predict the next token in a sequence, creating a robust framework for understanding and generating protein sequences. To address the challenges inherent in working with a limited training dataset, we implemented Low-Rank Adaptation (LoRA), an innovative fine-tuning technique. LoRA introduces trainable low-rank matrices into the model’s attention mechanisms, strategically reducing the number of trainable parameters while maintaining high performance. Our specific LoRA configuration included a low-rank dimension of 16, a scaling factor of 32, a dropout rate of 0.1, and a weight decay of 0.01, carefully calibrated to optimize the model’s learning capabilities. The model was configured with a maximum sequence length of 512 tokens, providing sufficient flexibility to handle diverse input sequence lengths efficiently. We utilized a tokenizer consistent with the model’s architecture, replacing padding tokens with end-of-sequence tokens to ensure seamless compatibility with the causal language modeling objectives. To mitigate potential overfitting and enhance performance on the limited dataset, we incorporated several advanced optimization techniques. Regularization strategies, including dropout and weight decay, were applied to prevent the model from becoming too specialized to the training data. An early stopping mechanism was implemented, monitoring performance on a validation subset and halting training when performance plateaued.
Additional training optimizations included a learning rate of $5\times 10^{-4}$ to encourage rapid convergence, a small per-device batch size of 4 to balance memory constraints and training dynamics, and 50 warmup steps to ensure stability during the initial training phases. The entire training process was conducted over 10 epochs, leveraging the Hugging Face Trainer API for streamlined implementation.
A sophisticated checkpoint system was employed to retain the best-performing model configurations, limiting storage requirements to the top two iterations. This approach ensures that we capture the most promising model states while maintaining computational efficiency.
By integrating these advanced techniques, we developed a robust and adaptable fine-tuning methodology that maximizes the potential of our limited dataset, creating a powerful computational tool for protein sequence analysis and generation.
@@ 3.3 Stage 3: Level 2 Fine-Tuning for Sequence-Property Associations
In this stage, the model undergoes Level 2 fine-tuning to establish robust associations between MaSp repeat sequences and their corresponding mechanical properties, such as toughness and extensibility. The fine-tuning objective serves a dual purpose: generating MaSp repeats with targeted mechanical properties and predicting mechanical properties from given MaSp repeat sequences.
This bidirectional approach enhances the model’s ability to understand and predict sequence-property relationships, significantly advancing its utility in the design of synthetic spider silk proteins.
Dataset: For level 2 fine-tuning, we assembled MaSp sequences along with their corresponding mechanical properties from 293 spider species in the Spider Silkome database. From the MaSp repeat dataset, we filtered sequences that had corresponding mechanical property annotations in the Spider Silkome dataset, yielding 592 MaSp sequences specifically designed for learning sequence-property correlations. Of these, 572 instances were used for training, while 20 were reserved as a test set.
The mechanical property set includes toughness, young’s modulus (E), tensile strength, and strain at break, along with their respective standard deviations. These eight values, used as conditioning features during fine-tuning and conditional generation, are represented as follows in equation 1:
@@
$\mathbf{P}=\left[\text{Toughness},E,\text{Strength},\text{Strain}\right],\quad\mathbf{SD}=\left[\sigma_{\text{Toughness}},\sigma_{E},\sigma_{\text{Strength}},\sigma_{\text{Strain}}\right]$ (1)
Since these properties span different value ranges, we applied Min-Max normalization to scale all values between 0 and 1 before training. This normalization facilitates faster convergence and simplifies value comparison in data analysis.
To facilitate bidirectional training, we curated the dataset to include a task-specific token that distinguishes between the forward and reverse tasks. This ensures the model effectively learns both sequence generation and property prediction within a unified framework. Model’s input-output relationships during the bidirectional training process:
  • Forward Task: Given a set of mechanical properties, the model generates a corresponding MaSp repeat sequence. Mentioned in equation 2
    $\text{Task Token}_{\text{EstimateProperty}}+\text{MaSp repeat}\rightarrow\text{Model}\rightarrow\text{Mechanical Properties}_{\text{8D}}$ (2)
  • Reverse Task: Given a MaSp repeat sequence, the model predicts its mechanical properties. Mentioned in equation 3
    $\text{Task Token}_{\text{GenerateSequence}}+\text{Mechanical Properties}_{\text{8D}}\rightarrow\text{Model}\rightarrow\text{MaSp repeat}$ (3)
@@
By explicitly incorporating task tokens, we enable the model to generalize across both tasks, reinforcing its ability to capture meaningful sequence-property relationships.
To optimize the model’s performance on the limited dataset of 572 records, we employed the LoRA (Low-Rank Adaptation) approach for fine-tuning. We maintained the same configuration used in our level 1 fine-tuning phase, utilizing a low-rank dimension of 16 and a scaling factor of 32. For optimization, the model was trained with a learning rate of $1\times 10^{-4}$, ensuring stable convergence while adapting to the limited data. We also adjusted the batch size to 8, optimizing memory usage while ensuring stable training dynamics. The model was trained for 5 epochs, ensuring adequate learning without overfitting.
To enhance generalization and avoid overfitting, regularization techniques such as dropout and weight decay were applied. A dropout rate of 0.2 and weight decay of 0.02 helped control model complexity. Additionally, an early stopping mechanism was implemented to halt training when performance on a validation subset plateaued, preventing unnecessary overfitting.
The refined model demonstrates exceptional capabilities in deciphering the complex relationships between protein sequences and their mechanical properties. By bridging computational modeling with experimental insights, this approach provides researchers and engineers with a sophisticated computational tool for designing synthetic spider silk proteins. The model’s enhanced predictive accuracy represents a significant advancement in biomaterial design, offering unprecedented insights into the intricate connections between molecular structure and material performance. This methodology not only addresses the challenges of working with limited experimental data but also establishes a robust framework for computational protein engineering. The approach showcases the potential of machine learning techniques to extract meaningful insights from complex biological systems, opening new avenues for targeted material design and scientific innovation.
@@ 4.3 Unconditional Sequence Generation: Self-consistency Assessment
To evaluate the model’s ability to generate biologically plausible sequences, we synthesized 200 sequences unconditionally, meaning no specific mechanical property constraints were applied (mentioned in supporting document SD1). Their alignment with natural MaSp sequences was analyzed by comparing key property distributions against 592 MaSp repeats used in level 2 fine-tuning. The results, illustrated in Figure 3, provide a quantitative measure of sequence fidelity, offering insight into how well the model captures the fundamental characteristics of spidroins.
@@
To further assess the biological plausibility of the generated sequences, we conducted a comprehensive analysis of key physiochemical attributes, including sequence length, molecular weight (Appendix A.2), instability index, and isoelectric point [44]. These properties were computed using ProtParam [45], a widely used tool for protein sequence analysis, implemented via a Python-based tool package to ensure accuracy and reproducibility. The results indicate that the generated sequences exhibit similar distributions across these physiochemical attributes, aligning closely with natural MaSp sequences.
Additionally, secondary structure fractions were analyzed using the ProteinAnalysis module from Biopython [46], which follows established secondary structure prediction frameworks, including DSSP [47] and Chou-Fasman propensity scales [48]. The ProteinAnalysis.secondary_structure_fraction() function was used to compute the fractional composition of $\alpha$-helices, strands and others/unstructured regions. This analysis enabled a quantitative comparison of secondary structure content between natural and generated sequences, providing insights into structural stability and folding tendencies.
Furthermore, the frequency of key motifs in the generated sequences was examined as part of the secondary structure evaluation. Figure 3 presents the distribution of poly-Ala, GGX, YGQGG, and SV motifs, offering insight into their role in sequential integrity. The formal definition of these motifs is provided in Equation 4. The selected set of motifs are known to impact mechanical properties of the silk fiber [15, 26] (refer section 2.1). The motif frequency distribution of natural and generated sequences, illustrated in Figure 3, suggests that the generated sequences reflect a similar distribution as natural sequences.
To further explore the structural properties of the generated sequences, we analyzed the amino acid composition and secondary structure fractions, which are critical determinants of protein function and mechanical behavior. Figures 3 present these comparisons in detail. It’s worth mentioning that the standard secondary structure prediction methods are optimized for globular proteins, making them less reliable for structural proteins like MaSp repeats [30].

### v2

@@ 3.1 Stage 1: Distillation of ProtGPT2 on Spider Sequences
ProtGPT2 (Ferruz et al., 2022)stands out among generative protein language models (PLMs) due to its specialized training and demonstrated ability to generate biologically plausible protein sequences. Unlike large-scale PLMs such as ESM-2 (Lin et al., 2023) and ProteinMPNN (Dauparas et al., 2022), which focus on structure prediction and sequence design constrained by existing protein scaffolds, ProtGPT2 is designed specifically for de novo protein generation. It has been trained on 10 million protein sequences from the UniRef50 dataset of the UniProt Knowledgebase (UniProtKB) (The UniProt Consortium, 2024) using a self-supervised learning approach. This extensive pre-training enables ProtGPT2 to predict the next amino acid in a sequence, effectively capturing the underlying "grammar" of protein structures.
@@
To address this challenge, we employ a multilevel strategy that leverages a pre-trained protein language model while optimizing it for our specific task. We begin with ProtGPT2. To enhance efficiency and adapt the model for MaSp repeat generation, we apply knowledge distillation (Hinton, 2015), creating a smaller, task-specific variant: SpiderGPT. This distilled model retains the essential knowledge of its teacher while significantly reducing model size and improving inference time, making it more practical for generating novel MaSp repeat sequences tailored to specific mechanical properties.
Dataset: SpiderGPT was trained on a curated dataset of approximately 100k protein sequences obtained from UniProtKB. These sequences were specifically selected based on their taxonomic classification within Araneae (spiders) and an annotation score greater than 1. (Details of the procedure are provided in Appendix A.1.) This focused dataset enables SpiderGPT to specialize in spider proteins while leveraging the broader knowledge distilled from the teacher model.
Training Process: The distillation process follows a standard teacher–student framework:
  • Teacher: ProtGPT2, a large pre-trained protein language model, serves as the source of knowledge.
  • Student: SpiderGPT, a lightweight model trained to replicate the teacher’s behavior while maintaining computational efficiency.
Distillation was guided by several critical hyperparameters. The temperature parameter $T=10$ was used to soften the teacher’s output probability distributions, enabling the student to learn finer inter-token dependencies. An interpolation coefficient $\alpha=0.1$ controlled the trade-off between soft targets from the teacher and hard targets from the original training data, promoting both generalization and retention of ground-truth patterns.
The SpiderGPT model was implemented with an embedding dimension of $n_{\text{embd}}=512$, six transformer layers ($n_{\text{layer}}=6$), and eight self-attention heads ($n_{\text{heads}}=8$). This reduced architecture was selected to balance representational capacity with training efficiency, allowing the student model to retain critical domain knowledge while being computationally tractable.
For effective knowledge transfer, both teacher and student models were trained on the same tokenized dataset. The teacher model generated soft labels—probability distributions over the vocabulary for each token in the input—which served as training targets for the student. Unlike hard labels, these soft targets convey distributional information about alternative token probabilities, allowing the student model to learn nuanced structural and contextual relationships embedded in the teacher’s output.
@@ 3.2 Stage 2: Level 1 Fine-Tuning on MaSp Repeat Regions
Dataset: For this study, we utilized the Spider Silkome dataset (Arakawa et al., 2022), a comprehensive resource cataloging silk gene sequences from 1,098 spider species and measuring mechanical, thermal, structural, and hydration properties for 446 species. It highlights the role of MaSp paralogs (MaSp1–MaSp3) in high-performance silk and identifies key amino acid motifs contributing to silk properties. This dataset serves as an open platform for advancing biomaterial research and innovation.
@@
For level 1 fine-tuning, we curated a dataset of MaSp sequences, resulting in 6,000 instances. To focus on the functional repeat regions, we removed the first 150 and last 115 amino acid residues, which correspond to the N-terminal and C-terminal domains, respectively (De Oliveira et al., 2024). The repeat sequences of the same MaSp type were then concatenated for each species to create a structured dataset.
Let this dataset be represented as a collection of unlabeled MaSp repeat sequences:
$\mathcal{D}_{1}=\left\{s_{i}\mid s_{i}\in\Sigma^{*},\;i=1,\ldots,6000\right\}$
where $\Sigma$ is the standard amino acid alphabet and $s_{i}$ denotes a variable-length MaSp repeat sequence over $\Sigma$. This dataset was used to train the model to learn token-level dependencies and contextual representations across natural MaSp repeats.
Training Process: The fine-tuning procedure adopted a causal language modeling (CLM) objective, which exploits the auto-regressive nature of the transformer architecture. This objective enables the model to predict each token in a sequence conditioned on its preceding tokens, providing a principled framework for learning sequential dependencies in protein sequences.
To address the challenges of fine-tuning on a limited dataset, we utilized Low-Rank Adaptation (LoRA), a parameter-efficient technique that introduces trainable low-rank matrices into the attention mechanisms of the transformer. Rather than updating all weights, LoRA injects additional parameters into selected layers, substantially reducing the number of trainable weights while preserving model expressiveness. Our implementation followed the standard configuration from Hugging Face’s PEFT library. Specifically, we used a LoRA rank of $r=16$, which defines the dimensionality of the inserted low-rank matrices. A higher rank allows for increased adaptation capacity at the cost of more parameters. The scaling factor was set to $\alpha=32$, amplifying the contribution of the low-rank component before integration into the base model. To regularize training, a dropout rate of 0.1 was applied within the LoRA layers. Biases in the original model layers were left unmodified (bias=none), and a weight decay coefficient of 0.01 was used to further constrain parameter growth during optimization.
The model was configured with a maximum sequence length of 512 tokens to accommodate variability in input length while ensuring consistent tensor shapes during batching. The tokenizer was adapted from the base model and configured to replace padding tokens with end-of-sequence tokens, ensuring compatibility with the CLM objective.
Training optimization included a learning rate of $5\times 10^{-4}$ to enable rapid convergence, a small per-device batch size of 4 to manage memory constraints, and 50 warm-up steps to stabilize learning in the initial epochs. Training proceeded for a maximum of 10 epochs, with early stopping enabled based on validation loss to prevent overfitting. The Hugging Face Trainer API was used for streamlined integration of training, logging, and checkpointing.
To manage storage and preserve high-performing checkpoints, we retained only the top two model states based on validation metrics. This strategy ensured the availability of optimal models for downstream use while maintaining computational efficiency.
By integrating LoRA with careful regularization, early stopping, and resource-aware optimization, we established a robust fine-tuning pipeline capable of leveraging limited protein sequence data to train powerful generative models. This methodology provides a flexible and scalable foundation for downstream applications in protein design and biomolecular sequence modeling.
@@ 3.3 Stage 3: Level 2 Fine-Tuning for Sequence-Property Associations
To enable effective learning of sequence–property relationships in the MaSp dataset, we conducted Level 2 fine-tuning using a parameter-efficient fine-tuning (PEFT) strategy based on Low-Rank Adaptation (LoRA). This phase was designed to balance model generalization and computational efficiency, while maintaining evaluation rigor through a custom cross-validation framework. The fine-tuning objective was bidirectional in nature: (i) to generate MaSp repeat sequences with specified mechanical properties, and (ii) to predict mechanical properties given an input MaSp repeat.
This dual-task setup improves the model’s ability to learn robust mappings between sequence and function, significantly enhancing its utility in the rational design of synthetic spider silk proteins.
Dataset:
For the second level of fine-tuning, we curated a dataset of MaSp (major ampullate spidroin) repeat sequences and their corresponding mechanical properties from the Spider Silkome database, covering 293 spider species. To ensure high-quality and consistent annotations, we filtered the dataset to include only sequences with complete mechanical property data and excluded repeated entries at the MaSp subtype level. This filtering process yielded a final set of 592 unique MaSp repeat sequences, suitable for modeling sequence–property relationships. Each data point comprises the amino acid sequence of a MaSp repeat, along with four mechanical properties: toughness, Young’s modulus ($E$), tensile strength, and strain at break, along with their respective standard deviations.
Let this dataset be represented as a set of labeled samples:
$\mathcal{D}_{2}=\left\{\left(s_{i},\mathbf{c}_{i},t_{i}\right)\mid i=1,\ldots,592\right\}$
where:
  • $s_{i}\in\Sigma^{*}$ is a MaSp repeat sequence over the amino acid alphabet $\Sigma$,
  • $\mathbf{c}_{i}\in[0,1]^{8}$ is the normalized conditioning vector, defined as:
@@
    $\mathbf{c}_{i}=\left[\text{Toughness}_{i},\sigma_{\text{Toughness},i},\text{Strength}_{i},\sigma_{\text{Strength},i},E_{i},\sigma_{E,i},\text{Strain}_{i},\sigma_{\text{Strain},i}\right]$
    Each property is paired with its corresponding standard deviation, and all values are scaled to the range $[0,1]$ via Min–Max normalization.
  • $t_{i}\in\{\texttt{<GenerateSequence>},\texttt{<EstimateProperty>}\}$ is the task token prepended to the model input to specify the prediction direction.
The model learns both directions through task conditioning: predicting properties from sequences and generating sequences from target property profiles, using the full 8-dimensional vector $(\mathbf{p}_{i},\bm{\sigma}_{i})$ as conditioning input or target as appropriate.
To standardize the feature space and facilitate model convergence, all property values and standard deviations were scaled to the $[0,1]$ range using Min–Max normalization.
To support bidirectional training, we introduced a task-specific token at the beginning of each input, explicitly indicating whether the model should perform sequence generation or property estimation. This ensures clear task conditioning within a unified modeling framework. The two tasks are defined as:
  • Forward Task (Property Estimation):
    $\texttt{<EstimateProperty>}\;\|\;s_{i}\rightarrow\text{Model}\rightarrow\mathbf{c}_{i}$
  • Reverse Task (Sequence Generation):
    $\texttt{<GenerateSequence>}\;\|\;\mathbf{c}_{i}\rightarrow\text{Model}\rightarrow s_{i}$
@@
To adapt the model for learning sequence–property relationships from limited annotated data, we reused the LoRA-based architecture introduced during Level 1 fine-tuning. This included injecting trainable low-rank matrices into the attention mechanisms and leveraging the same dropout, weight decay, and optimization settings to maintain regularization and efficiency.
The model was trained using a learning rate of $1\times 10^{-5}$, with a batch size of 8 per device. We used 50 warm-up steps to stabilize early optimization and trained for up to 10 epochs. An early stopping mechanism was activated based on validation loss to avoid overfitting. Tokenization was performed using a custom tokenizer derived from the base model, with all sequences padded and truncated to a fixed maximum length of 512 tokens. The end-of-sequence token was used for padding to ensure consistency with the model’s CLM objective.
Cross-Validation: To evaluate model generalization under multiple data partitions while maintaining high training efficiency, we implemented a custom cross-validation strategy. The dataset was first divided into 16 equal subsets, each containing 37 sequences. Rather than training on all 16 folds, we randomly selected 5 representative folds to serve as validation sets in independent training runs. In each run, one subset was held out for validation, while the remaining 555 sequences were used for training. This approach allows for robust estimation of performance variance while controlling for training resource demands.
Each run employed a causal language modeling objective, in which the model predicts the next token in a sequence given its preceding context. After each training cycle, the fine-tuned model and tokenizer were saved separately, and the corresponding validation set was exported for downstream analysis. This approach enabled isolated evaluation across folds and supported robust assessment of generalization performance.
The resulting models exhibit a strong ability to capture complex, non-linear relationships between MaSp repeat amino acid sequences and their associated mechanical properties. By leveraging advanced transformer architectures in combination with principled fine-tuning and rigorous cross-validation, this approach offers a promising computational framework for the rational design of synthetic spider silk proteins. It highlights the potential of machine learning to derive meaningful insights from limited experimental data and demonstrates a scalable, generalizable strategy for biomaterial discovery and engineering.
@@ 4.3 Unconditional Sequence Generation: Self-consistency Assessment
To assess the model’s capability to generate biologically plausible sequences, we unconditionally generated 100 MaSp repeats from each of the five independently trained models in Level 2 fine-tuning (a sample of 50 sequences is provided in Supporting Document SD1). This resulted in a total of 500 sequences generated without conditioning on mechanical property values. To evaluate their fidelity, we compared the distribution of key biochemical and structural properties against the 592 natural MaSp repeats used during Level 2 fine-tuning. The results, presented in Figure 3, quantitatively assess the similarity between generated and natural sequences, highlighting how well the model captures fundamental features of spidroin repeat regions.
@@
To further assess the biological plausibility of the generated sequences, we performed a comprehensive analysis of key physiochemical attributes, including sequence length, molecular weight, instability index, and isoelectric point (see Appendix A.2 for more details). These properties were computed using ProtParam (Gasteiger et al., 2005), a widely used tool for protein sequence analysis, implemented via a Python-based tool package to ensure accuracy and reproducibility. The results indicate that the generated sequences exhibit similar distributions across these physiochemical attributes, aligning closely with natural MaSp sequences.
Additionally, secondary structure fractions were analyzed using the ProteinAnalysis module from Biopython (Cock et al., 2009), which follows established secondary structure prediction frameworks, including DSSP (Kabsch & Sander, 1983) and Chou-Fasman propensity scales (Chou & Fasman, 1974). The ProteinAnalysis.secondary_structure_fraction() function was used to compute the fractional composition of $\alpha$-helices, strands and others/unstructured regions. This analysis enabled a quantitative comparison of secondary structure content between natural and generated sequences, providing insights into structural stability and folding tendencies.
Furthermore, the frequency of key motifs in the generated sequences was examined as part of the secondary structure evaluation. Figure 3 presents the distribution of poly-Ala, GGX, YGQGG, and SV motifs, offering insight into their role in sequential integrity. The formal definition of these motifs is provided in Equation 1. The selected set of motifs are known to impact mechanical properties of the silk fiber (Arakawa et al., 2022; Nakamura et al., 2024b) (refer section 2.1). The motif frequency distribution of natural and generated sequences, illustrated in Figure 3, suggests that the generated sequences reflect a similar distribution as natural sequences.
To further explore the structural properties of the generated sequences, we analyzed the amino acid composition and secondary structure fractions, which are critical determinants of protein function and mechanical behavior. Figures 3 present these comparisons in detail. It’s worth mentioning that the standard secondary structure prediction methods are optimized for globular proteins, making them less reliable for structural proteins like MaSp repeats (Jumper et al., 2021).